### v1

@@ Abstract
Large Language Models (LLMs) have rapidly advanced, with Gemini-3-Pro setting a new performance milestone. In this work, we explore collective intelligence as an alternative to monolithic scaling, and demonstrate that open-source LLMs’ collaboration can surpass Gemini-3-Pro. We first revisit LLM routing and aggregation at scale and identify three key bottlenecks: (1) current train-free routers are limited by a query-based paradigm focusing solely on textual similarity; (2) recent aggregation methods remain largely static, failing to select appropriate aggregators for different tasks; (3) the complementarity of routing and aggregation remains underutilized. To address these problems, we introduce JiSi, a novel framework designed to release the full potential of LLMs’ collaboration through three innovations: (1) Query-Response Mixed Routing capturing both semantic information and problem difficulty; (2) Support-Set-based Aggregator Selection jointly evaluating the aggregation and domain capacity of aggregators; (3) Adaptive Routing-Aggregation Switch dynamically leveraging the advantages of routing and aggregation. Comprehensive experiments on nine benchmarks demonstrate that JiSi can surpass Gemini-3-Pro with only 47% costs by orchestrating ten open-source LLMs, while outperforming mainstream baselines. It suggests that collective intelligence represents a novel path towards Artificial General Intelligence (AGI).
@@
Figure 1: The overall leaderboard of different LLMs and the proposed JiSi. By orchestrating open-source LLMs collectively, JiSi surpasses all monolithic LLMs, including leading closed-source LLMs such as Gemini-3-Pro, suggesting that collective intelligence represents a novel path towards AGI.
@@ 1 Introduction
The landscape of Artificial Intelligence (AI) has been fundamentally reshaped by Transformer-based Large Language Models (LLMs) (vaswani2017attention; brown2020language). By continuously scaling up model parameters and training data, LLMs have achieved unprecedented capabilities (kaplan2020scaling; comanici2025gemini25pushingfrontier; claude-sonnet-4; gpt-5; gpt-5.2; gemini-3-pro). Most notably, Gemini-3-Pro (gemini-3-pro) has established a significant performance milestone across different fields, including knowledge and reasoning, coding and engineering, general chat and factuality. The overwhelming success of Gemini-3-Pro further reinforces the prevailing paradigm of training increasingly powerful monolithic models. However, this naturally raises a fundamental question: is the indefinite scaling of a monolithic “super model” the only path towards Artificial General Intelligence (AGI)?
@@
Figure 2: Comprehensive performance comparison between the proposed JiSi and other LLMs. (a) and (b) show the fine-grained capability comparison with other open-source and closed-source LLMs on different datasets, respectively. JiSi achieves superior performance across multiple benchmarks.
In this work, we instead explore collective intelligence as an alternative path, demonstrating that the collaboration of individually weaker open-source LLMs can surpass leading closed-source models, including Gemini-3-Pro. Existing approaches to multi-LLM collaboration can be broadly divided into two fundamental categories. The first category is routing-based methods. Representative approaches include RouterDC (chen2024routerdc), EmbedLLM (zhuang2024embedllmlearningcompactrepresentations), GraphRouter (feng2025graphroutergraphbasedrouterllm), Avengers (zhang2025avengerssimplerecipeuniting), and retrieval-based routers (tang2025opensourcellmscollaborationbeats). These methods typically rely on query embeddings based similarity matching, or lightweight performance predictors to select the most suitable model for each input. The second category is aggregation-based methods. Typical techniques include mixture-of-agents (wang2024mixtureofagentsenhanceslargelanguage), self-consistency (chen2023universalselfconsistencylargelanguage), and majority voting (chen2024llmcallsneedscaling). These methods seek to surpass the limitations of individual models by synthesizing complementary information or consolidating consensus from multiple LLMs.
To achieve our goal, we first revisit routing and aggregation at scale, under both giant LLMs and challenging benchmarks. We find that: (1) current routing methods (shnitzer2023large; chen2024routerdc; zhang2025avengerssimplerecipeuniting; zhuang2024embedllmlearningcompactrepresentations) are predominantly training-free, typically projecting queries into a latent space for clustering or matching based model selection. However, pre-trained embedding models only capture text similarity and fail to perceive deep semantics or task difficulty, leading to suboptimal or mismatched selection. (2) Recent aggregation techniques (wang2024mixtureofagentsenhanceslargelanguage; li2025rethinking; li2025smoa) mainly exploit complementary information but are largely static, making it difficult to adaptively select aggregators with strong aggregation and domain-specific capabilities, which inevitably suffers from bottlenecks at critical responses generation. (3) the complementarity between routing and aggregation is often overlooked: routing offers stability but is constrained by single model, aggregation can exceed individual model limits yet is vulnerable to low-quality outputs and noise (li2025smoa; wolf2025your; li2025rethinking), while straightforward combination approaches struggle to effectively leverage their respective strengths.
@@
We conduct comprehensive experimental comparisons and analyses to assess the effectiveness of the proposed method JiSi. Specifically, we select ten open-source giant LLMs as candidate models and evaluate JiSi across nine challenging benchmarks. As shown in Fig. 1, the proposed JiSi surpasses all current state-of-the-art closed-source proprietary models. Compared to Gemini-3-Pro, it achieves an average performance gain of +1.15 and saves 53.23% cost. As shown in Fig. 2, JiSi can take specific advantages of candidate LLMs and break through their upper bound. In addition, JiSi remarkably outperforms all mainstream baselines. Furthermore, we verify that JiSi, as a concise training-free framework, exhibits consistent performance improvement with the introduction of new LLMs, which highlights its strong scalability. These results suggest that collective intelligence is a potential new path towards AGI.
@@ 2 Related works
Collective intelligence has emerged as a promising paradigm for pushing the intelligence upper bound. By orchestrating collaboration mechanisms across multiple models, collective intelligence aims to substantially enhance performance on complex tasks. This paper investigates collective intelligence in multi-LLM collaboration, which can be broadly divided into routing-based and aggregation-based methods.
Router-based Methods. These approaches exploit the domain-specific strengths of heterogeneous models by routing queries based on task requirements. For instance, ZOOTER (lu2024routing) proposed to distill signals from off-the-shelf reward models to train a router capable of precisely distributing queries to LLMs with corresponding expertise. RouterDC (chen2024routerdc) enhances routing accuracy through dual contrastive learning. GraphRouter (feng2025graphroutergraphbasedrouterllm) constructs a heterogeneous task-query-LLM graph, formulating model selection as a dynamic edge prediction problem via GNN. Similarly, MODEL-SAT (zhang2025capability) encodes candidate model performance on core tasks into capability representations, utilizing a lightweight LLM to predict the optimal candidate. Avengers (zhang2025avengerssimplerecipeuniting) constructs a training-free router by clustering the question embeddings as the proxies of different question categories, and utilizing the performance of the clusters as a model profile to estimate the LLM candidates.
Aggregation-based Methods. These methods focus on integration, synthesizing independent outputs from multiple LLMs to generate a superior response. Agent-Forest (li2024more) employs a majority voting mechanism to select the most frequent consensus among responses. The Mixture of Agents (MoA) framework (wang2024mixtureofagentsenhanceslargelanguage) introduces a layered architecture that leverages the collective strengths of LLMs through iterative collaboration, significantly improving generation quality. Building on this, SMoA (li2025smoa) incorporates response selection and early stopping mechanisms to sparsify information flows among agents, balancing performance with efficiency. Self-MoA (li2025rethinking) explores aggregating responses repeatedly sampled from a single high-performance agent.
[…]
@@ 3 JiSi Framework
Rethinking Routers. Current LLM routers typically adopt a query-based scheme to select an LLM. To evaluate the connection between the given query and candidate LLMs, they directly extract the query representation via pre-trained embedding models. The query representation is used for clustering to build LLM profiles (zhang2025avengerssimplerecipeuniting; chen2024routerdc) or training LLM embedding (zhuang2024embedllmlearningcompactrepresentations). However, because the pre-trained embedding models are trained for general-purpose text retrieval, the extracted embedding focuses on the query text similarity while ignoring the deep essential information beyond text overlap, such as deep semantics and question difficulty. For example, “Prove every large even number is a sum of two primes” and “Prove every large even number is a sum of two odds” almost have the same words, except for the last word. However, the first question is far more difficult than the second one. Based on query embeddings, they are hardly distinguished. It results in suboptimal query features or corresponding query support sets, which leads to routing these two questions to the same LLM incorrectly.
To address this issue, JiSi leverages the potent semantic extraction capabilities of LLMs to obtain auxiliary information from both generated responses and token costs. Specifically, the response to a query provides latent semantic information beyond surface-level text similarity. For instance, an LLM may generate significantly different responses to the two aforementioned questions, enabling the embedding model to effectively distinguish them. Moreover, based on an intuitive heuristic that a more difficult query typically correlates with more token costs for the same LLM, JiSi incorporates token costs to achieve a more precise calibration of query difficulty. Notably, query text similarity remains indispensable as it captures essential coarse-grained semantic features. Thus, by a weighted integration of the query, response, and token costs, JiSi achieves more accurate model routing.
@@
Rethinking Aggregation. Existing aggregation methods (li2025smoa; tang2025opensourcellmscollaborationbeats) are heuristic and statically select an aggregator based on the overall performance, which focuses on the comprehensive ability of aggregators. It may not be suitable for a specific task. However, while an LLM’s aggregation proficiency is inherently correlated with its comprehensive capability, a domain-specific expert does not necessarily possess superior aggregation skills. Consequently, the ideal aggregator must strike a balance between domain-specific and comprehensive ability. To identify the optimal task-specific aggregator, JiSi evaluates candidates by considering both their comprehensive performance across broader queries and their specialized expertise within the specific knowledge domain of the given query. It is worthnoting that since Fig. 3 aims to give an intuitive visualization of the proposed method, the practical deployment is different. JiSi utilizes the coarse-grained characteristic of query text similarity. It constructs a larger query support set in a query bank via query-based similarity that is rich in task information but contains diverse categories of queries, which guarantees the balanced estimation of domain-specific expertise and comprehensive ability. In this manner, JiSi can dynamically identify a domain-specific and synthesis-skilled LLM for a given new query.
Rethinking Combination. Existing methods (chen2025symbolic; tang2025opensourcellmscollaborationbeats) generally adopt an inflexible combination of routing and aggregation, typically following “routing many and aggregating all” workflow that treats aggregation as a post-processing step of routing. However, the complementarity and dynamics of routing and aggregation have long been overlooked. Routers may be advantageous for some simple and direct tasks, whereas aggregation methods can be useful for other complex tasks. Thus, always applying aggregation is not efficiency with a large range of query complexity. Moreover, while routers are more stable because it only selects high-quality LLMs, it hardly breaks through the upper bound of single-model performance; conversely, aggregation can exceed the capability of individual models but suffers from inferior or noisy inputs. To synergize the advantages of routing and aggregation, we propose an adaptive routing-aggregation switch mechanism for output strategies based on prior scores. Specifically, JiSi utilizes a pre-built embedding bank during the routing phase to obtain a precise prior score for each candidate response. It prompts the pruning of inferior or noisy responses and allows the system to adaptively switch between aggregation modes of varying input numbers. Notably, if only a single response exceeds the high-score threshold, the strategy bypasses the aggregation process and applies direct routing, significantly enhancing inference efficiency.
@@ 3.1 Framework Overview
In summary, JiSi adopts a minimalist “route-and-aggregate” topology: based on a pre-built embedding bank, it conducts a unified LLM routing and responses aggregation. For embedding bank construction, an extensible set of queries is sampled from heterogeneous sources. Responses from candidate LLMs are collected for these questions. Afterwards, all queries and responses are fed into a pre-trained embedding model to obtain latent representations, which form the embedding bank. Based on the embedding bank, the processing of a new query comprises three essential components. 1) Support-Set-based Aggregator Selection: A support query set is selected for aggregator dispatching and then filtered to improve the assignment of candidate aggregatee models. 2) Query-Response Mixed Routing: For each model selected from the filtered support set, a composite score is derived from query, response, and reasoning-length similarities, serving as a prior evaluation of the model’s expertise capacity for the given query. 3) Adaptive Routing-Aggregation Switch: the practical aggregatees are determined based on their scores and response length to filter inferior responses and noise. If the number of aggregatees exceeds 1, aggregation is performed to obtain the final response; otherwise, the final response is retrieved from the single aggregatee, i.e., it falls back to top-1 routing. An overview of the framework is illustrated in Fig. 4.
@@ 3.2 Embedding Bank
The embedding bank is constructed as an LLM profile storage, facilitating architecture-agnostic evaluation. Intuitively, the embedding bank is built on an extensive collection of queries from diverse sources. The original queries, LLM responses, and correctness information are subsequently projected into continuous embedding spaces to support similarity analysis.
Specifically, let $\left[U\right]=\left[1,U\right]\cap\mathbb{N}$. Assuming that $N$ questions $\mathcal{Q}=\left\{\left.x_{i}\in\mathcal{S}\right|i\in\left[N\right]\right\}$ each with corresponding label $y_{i}\in\mathcal{S}$, and $M$ LLMs $\mathcal{A}=\left\{\boldsymbol{A}_{1},\boldsymbol{A}_{2},\dots,\boldsymbol{A}_{M}\right\},\,\boldsymbol{A}_{i}:\mathcal{S}\to\mathcal{S}$ are predetermined. The $\mathcal{S}$ contains all literal strings. Responses of each LLM are collected over all questions, obtaining $\mathcal{R}=\left\{\left.\hat{y}_{i,j}=\boldsymbol{A}_{j}\left(x_{i}\right)\right|i\in\left[N\right],j\in\left[M\right]\right\}$ with reasoning costs (completion tokens) $c_{i,j}\in\mathbb{R}_{+}$. A pre-trained embedding model $\boldsymbol{A}_{emb}:\mathcal{S}\to\mathbb{R}^{d}$ is introduced to map questions and responses into contiguous latent space, forming the embedding bank $\mathcal{E}$:

### v2

@@ Abstract
Large Language Models (LLMs) have rapidly advanced, with Gemini-3-Pro setting a new performance milestone. In this work, we explore collective intelligence as an alternative to monolithic scaling, and demonstrate that open-source LLMs’ collaboration can surpass Gemini-3-Pro. We first revisit LLM routing and aggregation at scale and identify three key bottlenecks: (1) current train-free routers are limited by a query-based paradigm focusing solely on textual similarity; (2) recent aggregation methods remain largely static, failing to select appropriate aggregators for different tasks; (3) the complementarity of routing and aggregation remains underutilized. To address these problems, we introduce JiSi, a novel framework designed to release the full potential of LLMs’ collaboration through three innovations: (1) Query-Response Mixed Routing capturing both semantic information and problem difficulty; (2) Support-Set-based Aggregator Selection jointly evaluating the comprehensive and domain capacity of aggregators; (3) Adaptive Routing-Aggregation Switch dynamically leveraging the advantages of routing and aggregation. Comprehensive experiments on nine benchmarks demonstrate that JiSi can surpass Gemini-3-Pro with only 47% costs by orchestrating ten open-source LLMs, while outperforming mainstream baselines. It suggests that collective intelligence represents a novel path towards Artificial General Intelligence (AGI).
@@
thanks: https://github.com/magent4aci/openJiSi
Figure 1: By orchestrating open-source LLMs collectively, JiSi surpasses all monolithic LLMs, including leading closed-source LLMs such as Gemini-3-Pro, suggesting that collective intelligence represents a novel path towards AGI.
@@ 1 Introduction
The landscape of Artificial Intelligence (AI) has been fundamentally reshaped by Transformer-based Large Language Models (LLMs) (Vaswani et al., 2017; Brown et al., 2020). By continuously scaling up model parameters and training data, LLMs have achieved unprecedented capabilities (Kaplan et al., 2020; Comanici et al., 2025; Anthropic, 2025a; OpenAI, 2025b, a; Google, 2025). Most notably, Gemini-3-Pro (Google, 2025) has established a significant performance milestone across different fields, including knowledge and reasoning, coding and engineering, general chat and factuality. The overwhelming success of Gemini-3-Pro further reinforces the prevailing paradigm of training increasingly powerful monolithic models. However, this naturally raises a fundamental question: is the indefinite scaling of a monolithic “super model” the only path towards Artificial General Intelligence (AGI)?
@@
Figure 2: Comprehensive performance comparison between the proposed JiSi and other LLMs. (a) and (b) show the fine-grained capability comparison with open-source/closed-source LLMs on different datasets. JiSi achieves superior performance across multiple benchmarks.
In this work, we instead explore collective intelligence as an alternative path, demonstrating that the collaboration of individually weaker open-source LLMs can surpass leading closed-source models, including Gemini-3-Pro. To achieve this goal, we first revisit routing and aggregation at scale, including giant LLMs and challenging benchmarks. We find that: (1) Current routing methods (Shnitzer et al., 2023; Chen et al., 2024b; Zhang et al., 2025b) typically project queries into a latent space for clustering or matching-based model selection. However, pre-trained embedding models only capture text similarity and fail to perceive deep semantics or task difficulty, leading to suboptimal or mismatched selection. (2) Recent aggregation techniques (Wang et al., 2024a; Li et al., 2025b, a) mainly exploit complementary information but are largely static, making it difficult to adaptively select aggregators with strong aggregation and domain-specific capabilities, which inevitably suffer from bottlenecks in critical response generation. (3) The complementarity between routing and aggregation is overlooked: routing offers stability but is constrained by a single model, aggregation can exceed individual model limits yet is vulnerable to low-quality outputs and noise (Li et al., 2025a; Wolf et al., 2025; Li et al., 2025b).
@@
We conduct comprehensive experimental comparisons and analyses to assess the effectiveness of the proposed method JiSi. Specifically, we select ten open-source giant LLMs as candidate models and evaluate JiSi across nine challenging benchmarks. As shown in Fig. 1 and Table 2, the proposed JiSi surpasses all current state-of-the-art closed-source proprietary models. Compared to Gemini-3-Pro, it achieves an average performance gain of +1.15 and saves 53.23% cost. As shown in Fig. 2, JiSi can take specific advantages of candidate LLMs and break through their upper bound. In addition, JiSi remarkably outperforms all mainstream baselines. Furthermore, we verify that JiSi, as a concise training-free framework, exhibits consistent performance improvement with the introduction of new LLMs and tasks, which highlights its strong scalability. It suggests that collective intelligence is a potential new path towards AGI.
@@ 2 Related works
This paper investigates collective intelligence in multi-LLM collaboration, which can be broadly divided into router-based and aggregation-based methods.
Router-based Methods. RouterDC (Chen et al., 2024b) enhances routing accuracy through dual contrastive learning. GraphRouter (Feng et al., 2025) constructs a heterogeneous task-query-LLM graph, formulating model selection as a dynamic edge prediction problem via GNN. Similarly, MODEL-SAT (Zhang et al., 2025a) encodes candidate model performance on core tasks into capability representations, utilizing a lightweight LLM to predict the optimal candidate. Avengers (Zhang et al., 2025b) constructs a training-free router by clustering the question embeddings as the proxies of different question categories, and utilizing the performance of the clusters as a model profile to estimate the LLM candidates.
Aggregation-based Methods. Agent-Forest (Li et al., 2024a) employs a majority voting mechanism to select the most frequent consensus among responses. The Mixture of Agents (MoA) framework (Wang et al., 2024a) introduces a layered architecture that leverages the collective strengths of LLMs through iterative collaboration, significantly improving generation quality. Building on this, SMoA (Li et al., 2025a) incorporates response selection and early stopping mechanisms to sparsify information flows among agents, balancing performance with efficiency. Self-MoA (Li et al., 2025b) explores aggregating responses repeatedly sampled from a single high-performance agent.
Besides, recent works have explored unified frameworks that combine routing and aggregation. For instance, Sybomlic-MOE (Chen et al., 2025b) adopts a skill-aware dynamic expert selection strategy, matching tasks to experts based on validation-derived skill keywords and aggregating their outputs. SMACS (Tang et al., 2025) offers a scalable alternative by selecting candidate LLMs through a retrieval-based prior selection and refining outputs via an exploration–exploitation-driven posterior enhancement.
@@ 3 JiSi Framework
Rethinking Routers. Current LLM routers use pre-trained embedding models to extract query representations for selecting LLMs (Zhang et al., 2025b; Chen et al., 2024b; Zhuang et al., 2024). While these methods focus on text similarity, they often miss deeper semantics and question difficulty. For instance, the queries ”Prove every large even number is a sum of two primes” and ”Prove every large even number is a sum of two odds” are similar in wording, but the first is much more difficult. This oversight leads to misrouting of similar queries. To address this, JiSi incorporates latent semantics from generated responses and token costs. Query responses provide deeper insights beyond simple text similarity, enabling effective differentiation based on varying responses. JiSi also factors in token costs, assuming that harder queries incur higher costs, enabling more accurate calibration of query difficulty. By combining query, response, and token costs with appropriate weights, JiSi improves model routing. JiSi uses Query-Response Mixed Routing to infer implicit metrics from LLM responses, such as problem difficulty, to improve router decisions. The integration of the response information helps disambiguate semantically similar but divergent problems. The proposed Query-Response Mixed Routing can capture this difference due to the strong semantic understanding ability of LLMs and provide a better query support set for routing suitable LLMs.
@@
Rethinking Aggregation. Existing aggregation methods (Li et al., 2025a; Tang et al., 2025) statically select an aggregator based on overall performance, focusing on the comprehensive ability of aggregators, and are inappropriate for a specific task. On the other side, a domain-specific expert does not possess superior aggregation skills. Consequently, the ideal aggregator must strike a balance between domain-specific and comprehensive ability. To identify the optimal task-specific aggregator, JiSi evaluates candidates by considering both their overall performance across broader queries and their specialized expertise within the query’s knowledge domain. It is worth noting that, since Fig. 3 aims to provide an intuitive visualization of the proposed method, the practical deployment differs. JiSi utilizes the coarse-grained characteristic of query text similarity. It constructs a larger query support set in a query bank via query-based similarity, which is rich in task information and includes diverse query categories, thereby ensuring balanced estimation of domain-specific expertise and comprehensive ability. In this manner, JiSi can dynamically identify a domain-specific, synthesis-skilled LLM for a new query.
Rethinking Combination. Existing methods (Chen et al., 2025b; Tang et al., 2025) typically use a rigid ”routing many and aggregating all” approach, treating aggregation as a post-processing step. This overlooks the complementary roles of routing and aggregation; routers are effective for simple tasks while aggregation excels in complex scenarios. Constantly using aggregation can be inefficient due to varying query complexities. Routers select high-quality LLMs, but cannot exceed individual model performance, while aggregation struggles with noisy inputs. To address this, we propose an adaptive routing-aggregation switch mechanism that uses prebuilt embeddings to score candidate responses. This system can prune inferior responses and switch between aggregation and routing. If merely one response exceeds a high-score threshold, it is routed directly, improving inference efficiency.
@@ 3.1 Framework Overview
In summary, JiSi adopts a minimalist “route-and-aggregate” topology: Based on a pre-built embedding bank, it conducts a unified LLM routing and responses aggregation. For embedding bank construction, an extensible set of queries is sampled from heterogeneous sources. Responses from candidate LLMs are collected for these questions. Afterwards, all queries and responses are fed into a pre-trained embedding model to obtain latent representations, which form the embedding bank. Based on the embedding bank, The processing of a new query comprises three essential components. 1) Support-Set-based Aggregator Selection: A support query set is selected for aggregator dispatching and then filtered to improve the assignment of candidate aggregatee models. 2) Query-Response Mixed Routing: For each model selected from the filtered support set, a composite prior score is derived from query, response, and reasoning-length similarities, serving as a prior evaluation of the model’s expertise for the given query. 3) Adaptive Routing-Aggregation Switch: The practical aggregatees are determined based on their prior scores and response length to filter inferior responses and noise. If the number of aggregatees exceeds 1, aggregation is performed to obtain the final response; otherwise, the final response is retrieved from the single aggregatee, i.e., falls back to top-1 routing. An overview of the framework is illustrated in Fig. 4.
@@ 3.2 Embedding Bank
The embedding bank is constructed as an LLM profile storage, facilitating architecture-agnostic evaluation. Considering the inherent heterogeneity and black-box nature of candidate LLMs Intuitively, the embedding bank is built on an extensive collection of queries from diverse sources. The original queries, LLM responses, and correctness information are subsequently projected into continuous embedding spaces to support similarity analysis. Specifically, let $\left[U\right]=\left[1,U\right]\cap\mathbb{N}$. Assuming that $N$ questions $\mathcal{Q}=\left\{\left.x_{i}\in\mathcal{S}\right|i\in\left[N\right]\right\}$ each with corresponding label $y_{i}\in\mathcal{S}$, and $M$ LLMs $\mathcal{A}=\left\{\boldsymbol{A}_{1},\boldsymbol{A}_{2},\dots,\boldsymbol{A}_{M}\right\},\,\boldsymbol{A}_{i}:\mathcal{S}\to\mathcal{S}$ are predetermined. The $\mathcal{S}$ contains all literal strings. Responses of each LLM are collected over all questions, obtaining $\mathcal{R}=\left\{\left.\hat{y}_{i,j}=\boldsymbol{A}_{j}\left(x_{i}\right)\right|i\in\left[N\right],j\in\left[M\right]\right\}$ with reasoning costs (completion tokens) $c_{i,j}\in\mathbb{R}_{+}$. A pre-trained embedding model $\boldsymbol{A}_{emb}:\mathcal{S}\to\mathbb{R}^{d}$ is introduced to map questions and responses into contiguous latent space, forming the embedding bank $\mathcal{E}$: